# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

### 06 Mar 2026 v1

@@ Abstract
We prove that embedded infinite plane triangulations in ergodic scale-free environments are close to their circle packing and Riemann uniformization embedding on a large scale, as long as suitable moment and connectivity conditions are satisfied. Ergodic scale-free environments were earlier considered by Gwynne, Miller and Sheffield (2018) in the context of the invariance principles for random walk, and they arise naturally in the study of random planar maps and Liouville quantum gravity.
@@ 1 Introduction
In this paper, we work with random infinite triangulations under the two aforementioned conformal embeddings. Circle packing for random triangulations have been studied in e.g. [BS01, GN13, ABG+16, GJN19, IM23, BG24], and the Riemann uniformization embeddings for random planar maps have been considered in e.g. [GR13, CUR15]. Our main result is, for a family of infinite plane triangulations in ergodic scale-free environments, they are close to their circle packing and Riemann uniformization embedding on a large scale. Our result holds under a moment condition on the degrees of the triangulation and the regularity of the environment, along with a condition guaranteeing that the embedded triangulation is sufficiently well-connected.
@@
The rest of this section is organized as follows. In Section 1.1 we introduce some backgrounds on circle packings and the Riemann uniformization embedding. In Section 1.2 we introduce ergodic scale-free cell configurations. In Section 1.3 we state our main results, Theorems 1.7 and 1.8. In Section 1.4 we discuss some applications and future directions related to this work. Finally in Section 1.5, we give a detailed outline of our proof of Theorems 1.7 and 1.8 in the rest of the paper.
@@ 1.1 Discrete conformal embeddings
Let $\mathcal{G}$ be a locally finite planar (multi)graph with vertex set $\mathcal{V}\mathcal{G}$ and edge set $\mathcal{E}\mathcal{G}$. Multiple edges connecting a pair of vertices are allowed, while loops (i.e., edges starting and ending at the same vertex) are not allowed. We say that a disjoint collection of points $\{z_{v}:v\in\mathcal{V}\mathcal{G}\}\subseteq\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ and a collection of simple (undirected) curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{G}\}$ on $\widehat{\mathbb{C}}$ form a planar embedding of $\mathcal{G}$, if the endpoints of curves $\gamma_{e}$ for $e$ connecting vertices $v,v^{\prime}$ are at $z_{v}$ and $z_{v^{\prime}}$, and different curves $\gamma_{e}$ and $\gamma_{e^{\prime}}$ may only intersect each other at starting and ending points. A planar map $\mathcal{M}$ is an embedding of a planar graph $\mathcal{G}$ modulo orientation-preserving homeomorphisms of $\widehat{\mathbb{C}}$. A face is a connected component of the complement of the planar map. A triangular face is a face with three edges on its boundary. A (planar) triangulation is the union of the vertices and edges of a planar map $\mathcal{M}$, as well as (a subset of) the triangular faces of $\mathcal{M}$. We will work with infinite plane triangulations, where the underlying planar map is one-ended and every face has degree three and belongs to the triangulation, and disk triangulations, where the triangulation is finite and topologically homeomorphic to the unit disk. See Section 2.1 for more details.
A planar map $\mathcal{M}$ can be embedded in $\widehat{\mathbb{C}}$ (as described in the previous paragraph) in a number of different ways. In this paper, we focus on two natural embeddings: the circle packing and the Riemann uniformization embedding. These two embeddings are canonical examples of discrete conformal embeddings. This means in particular that embeddings of the same planar map in different domains induce (or is believed to induce) a map that approximates the conformal map between the two domains. For circle packing, this was conjectured by Thurston and solved in [RS87] in the uniformly bounded degree case [^1] The paper only considers the case of the hexagonal lattice, but it is commented in [HS96] that the proof in [RS87] also works for uniformly bounded degrees. and by [HS96] in full generality. For Riemann uniformization embedding, this follows from the definition.
@@
A second natural way to embed random planar maps is through the Riemann uniformization theorem. Roughly speaking, for a triangulation $\mathcal{T}$, one associate each face with an equilateral triangle and glue them isometrically along their edges to get a Riemann surface $M(\mathcal{T})$; see Section 2.1 for details. If $\mathcal{T}$ is an infinite plane triangulation, there are again two possible cases, where the Riemann surface $M(\mathcal{T})$ is parabolic (resp. hyperbolic) if it is conformally equivalent to the whole plane (resp. the unit disk). As commented in [GR13], similar to the circle packing case, this is equivalent to the recurrence/transience of the Brownian motion on the surface $M(\mathcal{T})$. See e.g. [GR13, CUR15] for some works in this direction.
@@ 1.2 Ergodic scale-free cell configurations
    A planar map $\mathcal{M}$, which we call the associated map of $\mathcal{H}$, with vertices $\{v_{H}:H\in\mathcal{H}\}$. We require that if $v_{H}$ and $v_{H^{\prime}}$ are connected by an edge of $\mathcal{M}$ then $H\neq H^{\prime}$ and $H\cap H^{\prime}\neq\emptyset$.
@@
We emphasize that, in the definition of cell configurations, it is allowed that two cells $H$ and $H^{\prime}$ have nonempty intersections while $H$ and $H^{\prime}$ are not connected by any edge in $\mathcal{E}\mathcal{H}$. See Figure 1 for an illustration. On the other hand, we will need the following property on the connectedness of $\mathcal{H}$.
@@ 1.3 Main results
In this subsection, we state our main results on the closeness between cell configurations and their discrete conformal embeddings. We work on the case where the associated map $\mathcal{M}$ of $\mathcal{H}$ is an infinite plane triangulation. We emphasize that the cells in $\mathcal{H}$ do not need to be simply connected, and different cells may cross each other; see Figure 1. However, we do need the following almost planarity assumption, which says that $\mathcal{H}$ is close to some planar embedding of $\mathcal{M}$. In the rest of the paper, without specific mentioning, we will always assume that all the conductances are equal to 1, i.e., $\mathfrak{c}(e)=1$ for every edge $e\in\mathcal{E}\mathcal{H}$.
@@
Definition 1.5.
Let $\mathcal{H}$ be a cell configuration as in Definition $1.$ We say $\mathcal{H}$ is almost planar, if there exists a collection of points $\{z_{H}:H\in\mathcal{H}\}$ and simple curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{H}\}$ in $\mathbb{C}$ forming a planar embedding of the associated map of $\mathcal{H}$, and almost surely as $r\to\infty$,
$\frac{1}{r}\max_{H\in\mathcal{H}(B(0;r))}\max_{e}\,\operatorname{dist}(\gamma_{e},H)\to 0,$ (1.2)
@@
Let $\mathcal{H}$ be a random cell configuration such that its associated map is a simple infinite plane triangulation. Suppose $\mathcal{H}$ satisfies all the conditions in Definitions 1.3, 1.4 and 1.5. Further assume $\mathcal{H}$ satisfies the moment bound
@@
Then $\mathcal{H}$ is a.s. circle packing parabolic, and there exists a circle packing $\mathcal{P}$ for $\mathcal{H}$ and a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ such that the following is true. For $H\in\mathcal{H}$, let $c(H)=\frac{1}{\mathrm{area}(H)}\int_{H}z\,dz$ be the Euclidean center of $H$, and $o(H)$ be the center of the disk for $H$ in $\mathcal{P}$. Then almost surely,
@@
Furthermore, if $\mathcal{H}$ satisfies the rotation invariance in Definition 1.6, then $\mathbf{A}$ can be taken to be the identity matrix.
@@
Let $\mathcal{H}$ be a random cell configuration such that its associated map is an infinite plane triangulation. Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3, 1.4 and 1.5. Further assume $\mathcal{H}$ satisfies the moment bound (1.3). Then the Riemann surface $M(\mathcal{H})$ formed by the associated map of $\mathcal{H}$ is a.s. parabolic, and there exists a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ and a conformal map $\varphi_{0}:M(\mathcal{H})\to\mathbb{C}$ for $\mathcal{H}$ such that the following is true. For $H\in\mathcal{H}$, let $c(H)=\int_{H}z\,dz$ be the Euclidean center of $H$, and $v_{H}$ be the vertex for $H$ in $M(\mathcal{H})$. Then almost surely,
@@
In Section 5.2, we will prove that Theorem 1.7 also holds for general planar maps (non-triangulations) for a particular choice of the circle-packing and under a moment condition, where we recall that circle packings are typically not unique for general planar maps. In Section 5.3, we will show that Theorem 1.8 also holds if $\mathcal{M}$ is a $p$-angulation for any $p\geq 3$, where we recall that a $p$-angulation is a planar map where all faces have degree pp.
We comment that the convergence results of random planar maps under discrete conformal embeddings in [GMS20, GMS21, HS23, BGS25] are all stated for finite planar maps, whereas Theorems 1.7 and 1.8 are stated for infinite plane triangulations. In our subsequent work, we will apply Theorems 1.7 and 1.8 to work on the convergence of random planar maps under circle packing and Riemann uniformization embedding in both finite volume and infinite volume setting. More precisely, we will use a version of Theorems 1.7 and 1.8 (Theorem 5.1) with a weaker version of the line connectivity in Definition 1.4, which will be proved in Section 5.1.
@@ 1.4 Applications and future directions
In this section, we discuss some applications and future directions for this work. We start with the following two simple examples, which provide applications of Theorems 1.7 and 1.8. We believe the results in these examples are new, except that a finite-volume version of Example 1.9 for circle packings was proven in [IM23].
@@
An important motivation this work is that we, in future work, will apply the results of this paper to planar maps arising in the study of Liouville quantum gravity. Another future direction is to generalize our results to planar maps which are not triangulations. Let us discuss the latter direction in more detail.
For infinite plane triangulations, the circle packing is unique. For general planar maps, the circle packing is not necessarily unique; however, one can get uniqueness if one imposes further conditions; see e.g. [BS93, SCH92, THU02]. There are several different kinds of conditions that one can impose. One of the simplest ways to define a unique circle packing for general simple planar maps is as follows. Let $\mathcal{M}$ be a one-ended 2-connected simple planar map with face set $\mathcal{F}\mathcal{M}$. For each face $f\in\mathcal{M}$, we add a vertex $v_{f}$ and add edges to connect $v_{f}$ to the vertices incident to $f$. This turns $\mathcal{M}$ into a simple infinite plane triangulation $\mathcal{T}_{\mathcal{M}}$. We can consider the circle packing for $\mathcal{T}_{\mathcal{M}}$ and remove the disks for all the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$. This defines a circle packing for $\mathcal{M}$ and it is the unique, up to Möbius transform, circle packing constructed in this way. In Section 5.2, we will prove a variant of Theorem 1.7 for the version of the circle packing defined here.
On the other hand, for the Riemann uniformization embedding one has existence and uniqueness of the embedding for any planar map. In Section 5.3, we will show that Theorem 1.8 still holds if we assume our planar map is a $p$-angulation for $p=4,5,6,\dots$.
Problem 1.11.
Prove Theorem 1.7 for other variants of circle packings for non-triangulations, and prove Theorem 1.8 for non-$p$-angulations.
Regarding the Riemann uniformization embedding, we expect that general planar maps without loops can be treated by slightly modifying the argument of the current paper, possibly under a suitable probabilistic bound on the face degrees. The key step would be to prove that Lemma 4.2 still holds in this context. Furthermore, we believe it is also possible to treat maps with loops; in this case, it might additionally be necessary to give a different definition of $\phi_{0}$ in Section 4.2.
@@ 1.5 Outline
In this section we provide a moderately detailed overview of the content of the rest of the paper, as well as the outline of the proof of Theorems 1.7 and 1.8.
In Section 2, we present various preliminaries. In Section 2.1, we review various properties of circle packings and the Riemann surface associated with triangulations. In Section 2.2, gather some results on the ergodic averages of the ergodic modulo scaling cell configuration $\mathcal{H}$ in the framework of [GMS22, Section 2.2]. We will repeatedly use the following two facts: the maximum diameter of the cells which intersect a large origin-containing square is typically of smaller order than its side length (Lemma 2.6), and the sum of $\mathrm{diam}(H)^{2}\deg(H)^{4}$ over all the cells $H$ intersecting a large square $S$ which is not too far from origin is bounded from above by a deterministic constant times the area of $S$ (Lemma 2.7).
In Section 3, we prove Theorem 1.7. In Section 3.4 we prove that if the random walk on two different embeddings of the same infinite planar map both converge to planar Brownian motion, then the two embeddings must be close. In order to see this, we consider the wrapping around events for the random walk, and use this to prove that under appropriate scaling, the natural homeomorphisms induced by the embeddings are tight on compact sets. Then under the image of the limiting homeomorphism, the trace of a 2D Brownian motion still has the law of the trace of a 2D Brownian motion. In Proposition 3.14 we prove that any limiting homeomorphism must be a linear transform. By the almost planarity assumption it follows that in order to conclude the proof, it is sufficient to prove that the random walk with Dubejko weights on the cell configuration $\mathcal{H}$ and on the circle packing $\mathcal{P}$ of $\mathcal{H}$ both converge to planar Brownian motion.
@@
It remains to show that the random walk with Dubejko weights on the circle packing $\mathcal{P}$ converges to Brownian motion, which is carried out in Section $3.$ Following work by Bou-Rabee and Gwynne [BG24] it is sufficient to check that up to scaling, there are no macroscopic disks in $\mathcal{P}$. In [GJN19], Gurel-Gurevich, Jerison and Nachmias proved that the maximal diameter of the disks in a circle packing can be controlled by the inverse of Cannon’s vertex extremal length [CAN94]. In Proposition 3.13, using this and the averaging over $\mathrm{diam}(H)^{2}$ from Lemma 2.7, we prove that for any number $q_{0}$, there are at most finitely many cells with associated vertex extremal length at most $q_{0}$, which eventually leads to the convergence of the random walk on $\mathcal{P}$.
@@
In Section 4.2 we construct a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$. Similarly as in  [GMS22, Section 2.3], we start from a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ which serves as the (non-injective) a priori embedding of $M(\mathcal{H})$. Much care needs to be taken with the exact definition of $\phi_{0}$ and we choose to define it by mapping each vertex of $M(\mathcal{H})$ to a uniform point in the corresponding cell of $\mathcal{H}$, and then extend to every equilateral triangle in $M(\mathcal{H})$ via linear interpolation. For each $m\in\mathbb{N}$ we let $\phi_{m}$ be equal to $\phi_{0}$ on the boundary of the $\phi_{0}^{-1}$ image of certain large squares, and harmonic elsewhere. The Dirichlet energy of the piecewise linear function $\phi_{0}$ can be controlled via the average in Lemma 2.7, and the Dirichlet energy of $\phi_{m}-\phi_{m^{\prime}}$ is small when $m,m^{\prime}$ is large (Lemma 4.19). By standard properties of harmonic functions, this implies that the $\phi_{m}$’s are Cauchy in probability in the local uniform topology, and the limit $\phi_{\infty}$ of $\phi_{m}$ exists.
@@
In Section 4.3, we prove that the harmonic function $\phi_{\infty}$ is close to $\phi_{0}$ on a large scale (Proposition 4.20). We first follow the same outline as [GMS22, Section 2.4] to prove that $\phi_{\infty}$ is close to $\phi_{0}$ on a set of line segments, where in our setting the line segments are contained in the image of a suitable subset of $M(\mathcal{H})$ under a conformal map. To prove the closeness of $\phi_{\infty}$ to $\phi_{0}$ on the complementary set, the oscillation of $\phi_{\infty}$ in each connected component of the complement is equal to the oscillation of $\phi_{\infty}$ on the boundary of that component thanks to the maximum principle for harmonic functions, whereas the oscillation of $\phi_{0}$ in each connected component can be bounded by the continuity from Lemma $4.$ This shows that $|\phi_{0}-\phi_{\infty}|$ is also small in every connected component of the complementary set, which concludes the proof of Proposition 4.20.

### 06 May 2026 v2

# Circle packing and Riemann uniformization of random planar maps in an ergodic scale-free environment

@@ Abstract
We prove that embedded infinite planar maps in ergodic scale-free environments are close to their circle packing and Riemann uniformization embedding on a large scale, as long as suitable moment and connectivity conditions are satisfied. Ergodic scale-free environments were earlier considered by Gwynne, Miller and Sheffield (2018) in the context of the invariance principles for random walk, and they arise naturally in the study of random planar maps and Liouville quantum gravity.
@@ 1 Introduction
In this paper, we work with random planar maps with whole-plane topology, i.e., one-ended planar maps without boundary, under the two aforementioned conformal embeddings. Circle packing for random triangulations have been studied in e.g. [BS01, GN13, ABG+16, GJN19, IM23, BG24], and the Riemann uniformization embeddings for random planar maps have been considered in e.g. [GR13, CUR15]. Our main result is, for a family of random planar maps with whole-plane topology in ergodic scale-free environments, they are close to their circle packing and Riemann uniformization embedding on a large scale. Our result holds under a moment condition on the degrees of the triangulation and the regularity of the environment, along with a condition guaranteeing that the embedded triangulation is sufficiently well-connected.
@@
The rest of this section is organized as follows. In Section 1.1 we introduce some backgrounds on circle packings and the Riemann uniformization embedding. In Section 1.2 we introduce ergodic scale-free cell configurations. In Section 1.3 we state our main results, Theorems 1.6 and 1.7. In Section 1.4 we discuss some applications and future directions related to this work. Finally in Section 1.5, we give a detailed outline of our proof of Theorems 1.6 and 1.7 in the rest of the paper.
@@ 1.1 Discrete conformal embeddings
Let $\mathcal{G}$ be a locally finite planar (multi)graph with vertex set $\mathcal{V}\mathcal{G}$ and edge set $\mathcal{E}\mathcal{G}$. Multiple edges connecting a pair of vertices are allowed, while loops (i.e., edges starting and ending at the same vertex) are not allowed. We say that a disjoint collection of points $\{z_{v}:v\in\mathcal{V}\mathcal{G}\}\subseteq\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ and a collection of simple (undirected) curves $\{\gamma_{e}:e\in\mathcal{E}\mathcal{G}\}$ on $\widehat{\mathbb{C}}$ form a planar embedding of $\mathcal{G}$, if the endpoints of curves $\gamma_{e}$ for $e$ connecting vertices $v,v^{\prime}$ are at $z_{v}$ and $z_{v^{\prime}}$, and different curves $\gamma_{e}$ and $\gamma_{e^{\prime}}$ may only intersect each other at starting and ending points. A planar map $\mathcal{M}$ is an embedding of a planar graph $\mathcal{G}$ modulo orientation-preserving homeomorphisms of $\widehat{\mathbb{C}}$. The face set $\mathcal{F}\mathcal{M}$ is the set of the connected components of the complement of the planar map, and we say that $\mathcal{M}$ has whole-plane topology if all the faces are bounded in $\mathbb{C}$, and the union of faces in $\mathcal{F}\mathcal{M}$ along with their closures is equal to $\mathbb{C}$. A triangular face is a face with three edges on its boundary. A (planar) triangulation is the union of the vertices and edges of a planar map $\mathcal{M}$, as well as (a subset of) the triangular faces of $\mathcal{M}$. We will work with infinite plane triangulations, where the underlying planar map is one-ended and every face has degree three and belongs to the triangulation, and disk triangulations, where the triangulation is finite and topologically homeomorphic to the unit disk. See Section 2.1 for more details.
A planar map $\mathcal{M}$ can be embedded in $\widehat{\mathbb{C}}$ (as described in the previous paragraph) in a number of different ways. In this paper, we focus on two natural embeddings: the circle packing and the Riemann uniformization embedding. These two embeddings are canonical examples of discrete conformal embeddings. This means in particular that embeddings of the same planar map in different domains induce (or is believed to induce) a map that approximates the conformal map between the two domains. For circle packing, this was conjectured by Thurston and solved in [RS87] in the case of the hexagonal lattice, in [STE90, STE96] for circle packings with uniformly bounded radii ratio for neighboring disks via probabilistic methods, and by [HS96] in full generality. For Riemann uniformization embedding, this follows from the definition.
@@
A second natural way to embed random planar maps is through the Riemann uniformization theorem. Roughly speaking, for a planar map $\mathcal{M}$, one associate each face of degree $p$ with a regular $p$-gon of unit side length and glue them isometrically along their edges to get a Riemann surface $M(\mathcal{M})$; see Section 2.1 for details. If $\mathcal{M}$ has whole-plane topology, there are again two possible cases, where the Riemann surface $M(\mathcal{M})$ is parabolic (resp. hyperbolic) if it is conformally equivalent to the whole plane (resp. the unit disk). As commented in [GR13, AHN+18], similar to the circle packing case, this is equivalent to the recurrence/transience of the Brownian motion on the surface $M(\mathcal{M})$. See e.g. [GR13, CUR15, AHN+18] for some works in this direction.
@@ 1.2 Ergodic scale-free cell configurations
    A planar map $\mathcal{M}$, which we call the associated map of $\mathcal{H}$, with vertices $\{v_{H}:H\in\mathcal{H}\}$. We require that if $v_{H}$ and $v_{H^{\prime}}$ are connected by an edge of $\mathcal{M}$ then $H\cap H^{\prime}\neq\emptyset$.
@@
We emphasize that, in the definition of cell configurations, it is allowed that two cells $H$ and $H^{\prime}$ have nonempty intersections while $H$ and $H^{\prime}$ are not connected by any edge in $\mathcal{E}\mathcal{H}$. We also allow that the cells in $\mathcal{H}$ are not simply connected and may cross each other. See Figure 1 for an illustration. On the other hand, we will need the following property on the connectedness of $\mathcal{H}$.
@@ 1.3 Main results
In this subsection, we state our main results on the closeness between cell configurations and their discrete conformal embeddings. We start with the case where the associated map $\mathcal{M}$ of $\mathcal{H}$ is an infinite plane triangulation.
@@
Let $\mathcal{H}$ be a random cell configuration such that its associated map is a simple infinite plane triangulation. Suppose $\mathcal{H}$ satisfies all the conditions in Definitions 1.3 and 1.4. Further assume $\mathcal{H}$ satisfies the moment bound
@@
Then $\mathcal{H}$ is a.s. circle packing parabolic, and there exists a circle packing $\mathcal{P}$ for $\mathcal{H}$ and a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ such that the following is true. Almost surely, for any $\varepsilon>0$, there exists $R_{0}>0$ such that for every $r>R_{0}$, the maximal diameter of disks in $\mathcal{P}$ intersecting $B(0;r)$ is less than $\varepsilon r$. Furthermore, for $H\in\mathcal{H}$, letting $c(H)=\frac{1}{\mathrm{area}(H)}\int_{H}z\,dz$ be the Euclidean center of $H$, and letting $o(H)$ be the center of the disk for $H$ in $\mathcal{P}$, almost surely,
@@
Finally, if $\mathcal{H}$ satisfies the rotation invariance in Definition 1.5, then $\mathbf{A}$ can be taken to be the identity matrix.
@@
Let $\mathcal{H}$ be a random cell configuration such that its associated map is an infinite plane triangulation. Suppose $\mathcal{H}$ satisfies the conditions in Definitions 1.3 and 1.4. Further assume $\mathcal{H}$ satisfies the moment bound (1.2). Then the Riemann surface $M(\mathcal{H})$ formed by the associated map of $\mathcal{H}$ is a.s. parabolic, and there exists a deterministic $2\times 2$ matrix $\mathbf{A}$ with $\det(\mathbf{A})=1$ and a conformal map $\varphi_{0}:M(\mathcal{H})\to\mathbb{C}$ for $\mathcal{H}$ such that the following is true. For $H\in\mathcal{H}$, let $c(H)=\int_{H}z\,dz$ be the Euclidean center of $H$, and $v_{H}$ be the vertex for $H$ in $M(\mathcal{H})$. Then almost surely,
@@
For infinite plane triangulations, the circle packing is unique. For general planar maps, the circle packing is not necessarily unique; however, one can get uniqueness if one imposes further conditions; see e.g. [BS93, SCH92, THU02]. There are several kinds of conditions that one can impose. One of the simplest ways to define a unique circle packing for general simple planar maps is as follows. Let $\mathcal{M}$ be a 2-connected simple planar map with whole-plane topology and face set $\mathcal{F}\mathcal{M}$. For each face $f\in\mathcal{M}$, we add a vertex $v_{f}$ and add edges to connect $v_{f}$ to the vertices incident to $f$. This turns $\mathcal{M}$ into a simple infinite plane triangulation $\mathcal{T}_{\mathcal{M}}$. We can consider the circle packing for $\mathcal{T}_{\mathcal{M}}$ and remove the disks for all the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$. This defines a circle packing for $\mathcal{M}$ and it is the unique, up to Möbius transform, circle packing constructed in this way.
Let $\mathcal{H}$ be a cell configuration such that the associated map $\mathcal{M}$ is 2-connected and has whole-plane topology. Let $\mathcal{T}_{\mathcal{M}}$ be the triangulation generated from $\mathcal{M}$ as above. We write $v_{H}\sim f$ if the vertex $v_{H}$ for $H$ in the associated map $\mathcal{M}$ is incident to the face $f\in\mathcal{F}\mathcal{M}$. Let
be the union of cells whose vertex is on the same face as $v_{H_{0}}$ and let $d_{0}$ be the maximal degree of the faces incident to $v_{H_{0}}$.
Theorem 1.8.
Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ being 2-connected and having whole-plane topology. Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3 and 1.4. Further assume
Then the conclusions in Theorem 1.6 hold for the circle packing of $\mathcal{M}$ defined as above.
In the setting of the Riemann uniformization, let $\mathcal{M}$ be a planar map with whole-plane topology and face set $\mathcal{F}\mathcal{M}$. For each face $f$ enclosed by a single loop, we simply remove this loop. For each face $f\in\mathcal{F}\mathcal{M}$ with 2 edges, we contract the two edges to turn it into a single edge, such that each face of the resulting map has at least 3 edges. Let $M(\mathcal{M})$ be the Riemann surface obtained by associating each remaining face of degree $p\geq 3$ with regular $p$-gon of unit side length and glue them isometrically along their edges.
Theorem 1.9.
Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ having whole-plane topology. Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3 and 1.4. Further assume that (1.7) holds. Then the conclusions in Theorem 1.7 hold for the Riemann surface $M(\mathcal{M})$ defined as above.
As discussed in [AHN+18, Section 6], there are several other natural ways to define Riemann surfaces associated with planar maps; for instance, one can assign each face of $\mathcal{M}$ of degree $p$ a disk of circumference $p$, with boundary split into $p$ arcs length one corresponding to the edges, and glue adjacent faces according to arc length along their shared edges. We expect that similar conclusions still hold by minor modifications of our proof of Theorem 1.9.
We comment that the convergence results of random planar maps under discrete conformal embeddings in [GMS20, GMS21, HS23, BGS25] are all stated for finite planar maps, whereas Theorems 1.6-1.9 are stated for infinite plane triangulations. In our subsequent work, we will apply Theorems 1.6-1.9 to work on the convergence of random planar maps under circle packing and Riemann uniformization embedding in both finite volume and infinite volume setting. More precisely, we will use a version of Theorems 1.6-1.9 (Theorem 5.5) with a weaker version of the line connectivity in Definition 1.4, which will be proved in Section 5.3.
@@ 1.4 Applications and future directions
In this section, we discuss some applications and future directions for this work. We start with the following two simple examples, which provide applications of Theorems 1.6 and 1.7. We believe the results in these examples are new, except that a finite-volume version of Example 1.10 for circle packings was proven in [IM23].
@@
An important motivation this work is that we, in future work, will apply the results of this paper to planar maps arising in the study of Liouville quantum gravity.
@@ 1.5 Outline
In this section we provide a moderately detailed overview of the content of the rest of the paper, as well as the outline of the proof of Theorems 1.6 and 1.7.
In Section 2, we present various preliminaries. In Section 2.1, we review various properties of circle packings and the Riemann surface associated with planar maps. In Section 2.2, we gather some results on the ergodic averages of the ergodic modulo scaling cell configuration $\mathcal{H}$ in the framework of [GMS22, Section 2.2]. We will repeatedly use the following two facts: the maximum diameter of the cells which intersect a large origin-containing square is typically of smaller order than its side length (Lemma 2.6), and the sum of $\mathrm{diam}(H)^{2}\deg(H)^{4}$ over all the cells $H$ intersecting a large square $S$ which is not too far from origin is bounded from above by a deterministic constant times the area of $S$ (Lemma 2.7). As a consequence, similarly as in  [GMS22, Section 2.3], we start from a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ which serves as the (non-injective) a priori embedding of $M(\mathcal{H})$ and prove some of its topological properties. Much care needs to be taken with the exact definition of $\phi_{0}$ and we choose to define it by mapping each vertex of $M(\mathcal{H})$ to a uniform point in the corresponding cell of $\mathcal{H}$, and then extend to every equilateral triangle in $M(\mathcal{H})$ via linear interpolation. In Section 2.3, we recap the convergence of random walk on $\mathcal{H}$ to Brownian motion from [GMS22], and use it to prove that for each large square $S$, there is some planar embedding of $\mathcal{H}(S)$ that is close to $\mathcal{H}(S)$ (Proposition 2.15).
In Section 3, we prove Theorem 1.6. In Section 3.4 we prove that if the random walk on two different embeddings of the same infinite planar map both converge to planar Brownian motion, then the two embeddings must be close. In order to see this, we consider the wrapping around events for the random walk, and use this to prove that under appropriate scaling, the natural homeomorphisms induced by the embeddings are tight on compact sets. Then under the image of the limiting homeomorphism, the trace of a 2D Brownian motion still has the law of the trace of a 2D Brownian motion. In Proposition 3.12 we prove that any limiting homeomorphism must be a linear transform. By the almost planarity result from Proposition 2.15 it follows that in order to conclude the proof, it is sufficient to prove that the random walk with Dubejko weights on the cell configuration $\mathcal{H}$ and on the circle packing $\mathcal{P}$ of $\mathcal{H}$ both converge to planar Brownian motion.
@@
It remains to show that the random walk with Dubejko weights on the circle packing $\mathcal{P}$ converges to Brownian motion, which is carried out in Section 3.3. Following work by Bou-Rabee and Gwynne [BG24] it is sufficient to check that up to scaling, there are no macroscopic disks in $\mathcal{P}$. In [GJN19], Gurel-Gurevich, Jerison and Nachmias proved that the maximal diameter of the disks in a circle packing can be controlled by the inverse of Cannon’s vertex extremal length [CAN94]. In Proposition 3.11, using this and the averaging over $\mathrm{diam}(H)^{2}$ from Lemma 2.7, we prove that for any number $q_{0}$, there are at most finitely many cells with associated vertex extremal length at most $q_{0}$, which eventually leads to the convergence of the random walk on $\mathcal{P}$.
@@
In Section 4.2 we construct a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ using the function $\phi_{0}$ defined in Section 2.2. For each $m\in\mathbb{N}$ we let $\phi_{m}$ be equal to $\phi_{0}$ on the boundary of the $\phi_{0}^{-1}$ image of certain large squares, and harmonic elsewhere. The Dirichlet energy of the piecewise linear function $\phi_{0}$ can be controlled via the average in Lemma 2.7, and the Dirichlet energy of $\phi_{m}-\phi_{m^{\prime}}$ is small when $m,m^{\prime}$ is large (Lemma 4.17). By standard properties of harmonic functions, this implies that the $\phi_{m}$’s are Cauchy in probability in the local uniform topology, and the limit $\phi_{\infty}$ of $\phi_{m}$ exists.
@@
In Section 4.3, we prove that the harmonic function $\phi_{\infty}$ is close to $\phi_{0}$ on a large scale (Proposition 4.18). We first follow the same outline as [GMS22, Section 2.4] to prove that $\phi_{\infty}$ is close to $\phi_{0}$ on a set of line segments, where in our setting the line segments are contained in the image of a suitable subset of $M(\mathcal{H})$ under a conformal map. To prove the closeness of $\phi_{\infty}$ to $\phi_{0}$ on the complementary set, the oscillation of $\phi_{\infty}$ in each connected component of the complement is equal to the oscillation of $\phi_{\infty}$ on the boundary of that component thanks to the maximum principle for harmonic functions, whereas the oscillation of $\phi_{0}$ in each connected component can be bounded by the continuity from Lemma 4.3. This shows that $|\phi_{0}-\phi_{\infty}|$ is also small in every connected component of the complementary set, which concludes the proof of Proposition 4.18.